OpenStudy (anonymous):

the real fourth root of -81?

OpenStudy (experimentx):

x^4 = -81??

OpenStudy (anonymous):

I think he means ... \[\huge \sqrt[4]{-81}\]

OpenStudy (anonymous):

I don't know that it could be REAL ...

OpenStudy (mertsj):

-81 has no real fourth root

OpenStudy (experimentx):

no thats no way real.

OpenStudy (mertsj):

Perhaps you meant \[ x^4=81\]

OpenStudy (anonymous):

@keiko are you there? ... say something , how do you mean ?

OpenStudy (mertsj):

in which case, x=3 or -3